A 10kg fish swimming at 2.20m/sec suddenly gobbles up a 5kg fish that is initially stationary. Find the speed of the large fish just after it eats the small one.

Grade:6

1 Answers

Mukesh Sharma
10 years ago
Conservation of mometum problem
m1*v1i + m2*v2i = m1*v1f + m2*v2f
In this case after the smaller fish is eaten, they both move at the same speed
v1f = v2f = vf
and we know v2i = 0
m1*v1i = vf*(m1+m2)
vf = m1*v1i / (m1+m2)
vf = (10 kg * 2.20 m/s) / (10 kg + 5 kg)
vf =1.46 m/s
Thanks & Regards
Mukesh Sharma
askIITians Faculty